Make-or-Buy, Traditional Analysis

Wehner Company is currently manufacturing Part ABS-43, producing 55,900 units annually. The part is used in the production of several products made by Wehner. The cost per unit for ABS-43 is as follows:

Direct materials $46.35
Direct labor 10.60
Variable overhead 3.25
Fixed overhead 3.10
  Total $63.30

Of the total fixed overhead assigned to ABS-43, $14,702 is direct fixed overhead (the annual lease cost of machinery used to manufacture Part ABS-43), and the remainder is common fixed overhead. An outside supplier has offered to sell the part to Wehner for $60.06. There is no alternative use for the facilities currently used to produce the part. No significant non-unit-based overhead costs are incurred.

Required:

1. Should Wehner Company make or buy Part ABS-43?
Wehner should   the part. This will produce total cost savings of $.

2. What is the maximum amount per unit that Wehner would be willing to pay to an outside supplier? Round your answer to the nearest cent.
$ per unit
Answer #1

1. Total Relevant cost of Manufacturing = (Direct material + Direct Labor + Variable Overhead)*No of Unit Produced + direct fixed overhead

Total Relevant cost of Manufacturing = (46.35+ 10.60+ 3.25)*55900 + $14,702

Total Relevant cost of Manufacturing = $3,379,882

Total Cost of Buying = 55900*60.06

Total Cost of Buying = $3,357,354

Total cost savings if part is bought = 3,379,882 - 3,357,354

Total cost savings if part is bought = $22,528

Decision : Wehner should buy the part . This will produce total cost savings of $ 23,650.

2. Maximum amount per unit = Total Relevant cost of Manufacturing / No. of units = $3,379,882 / 55900 = $60.46 per unit